# Pulsar Timing Arrays require hierarchical models

### v1

@@ Abstract
Pulsar Timing Array projects have found evidence of a stochastic background of gravitational waves (GWB) using data from an ensemble of pulsars. In the literature, minimal assumptions are made about the signal and noise processes that affect data from these pulsars, such as pulsar spin noise. These assumptions are encoded as uninformative priors in Bayesian searches, though Frequentist approaches make similar assumptions. Uninformative priors are not suitable for (noise) properties of pulsars in an ensemble, and they bias estimates of model parameters such as gravitational-wave signal parameters. Both Frequentist and Bayesian searches are affected. In this letter, more appropriate priors are proposed in the language of Hierarchical Bayesian Modeling, where the properties of the ensemble of pulsars are jointly described with the properties of the individual components of the ensemble. Results by Pulsar Timing Array projects should be re-evaluated using Hierarchical Models.
@@ 1 Introduction
Astronomy is on the cusp of a transformative era with the advent of gravitational-wave (GW) astronomy. The LIGO-Virgo collaboration has heralded this new age with observations in the higher frequency bands (LIGO Scientific Collaboration and Virgo Collaboration et al., 2016), and now, Pulsar Timing Arrays (PTAs) stand poised to complement these findings in the nanohertz regime with published evidence for a stochastic background of gravitational waves (GWB) (Agazie et al., 2023a; Reardon et al., 2023; Antoniadis et al., 2023; Xu et al., 2023). This advancement holds the promise of reshaping our understanding of observational cosmology, galaxy formation, and physics of the early universe.
@@
PTAs represent a unique approach to gravitational-wave detection. By precisely monitoring the arrival times of pulses from an array of millisecond pulsars (MSPs, Backer et al., 1982), these projects seek to detect the subtle perturbations caused by passing gravitational waves (Sazhin, 1978; Detweiler, 1979; Romani, 1989; Foster & Backer, 1990). The stability of MSP emissions is key to this method; however, this stability is not without some level of irregularity. The rotational instabilities of the MSPs, often referred to as spin noise, presents a notable challenge for the PTA projects in their efforts to detect gravitational waves (Cordes & Shannon, 2010; Lentati et al., 2016).
@@
The complexity of analyzing an ensemble of pulsars simultaneously necessitates careful statistical modeling (Taylor, 2021), particularly in the treatment of spin noise. The PTA community has approached this with Bayesian inference, where separate spin noise components with uninformative amplitude parameters for each MSP are added to the model. While this practice seems intuitive, it has sometimes led to posterior bias in gravitational wave analyses, where I use the term posterior bias to describe the apparent shift of the posterior distribution with respect to the true model parameter values. Posterior bias can be a consequence of model misspecification. Multiple papers have pointed out that posterior bias occurs in contemporary analyses (Hazboun et al., 2020; Goncharov et al., 2022; Zic et al., 2022; Johnson et al., 2022), but no exact cause or satisfactory solution was given. These warning signals have been mostly ignored by subsequent searches for GWs. Those same priors are still used today, such as in recent high-profile publications (Agazie et al., 2023a; Reardon et al., 2023; Antoniadis et al., 2023; Xu et al., 2023).
@@
Even though currently used uninformative priors have been confirmed to work with injection recoveries in the literature (e.g. Lentati et al., 2013; van Haasteren et al., 2009; Ellis et al., 2013; Johnson et al., 2023, and most other methods papers on GWB detection), no rigorous statistical assessment of (posterior) bias has been possible. What has been done are studies that focus on the validity of the posterior distribution and its implementation, which can be done with $p$-$p$ plots of ensembles of mock data (e.g. Ellis et al., 2013). However, for those $p$-$p$ plots to be consistent, one needs to use the same priors for data generation as for analysis. By doing that, one glosses over a significant source of posterior bias: the use of uninformative priors on an ensemble of observable quantities that follow a different distribution in reality than assumed in the prior (Priestley, 1965).
@@
Frequentist estimators are similarly affected. A minimum variance unbiased estimator (MVUE) (Anholm et al., 2009; Chamberlin et al., 2015; Allen & Romano, 2023) needs to assume the data distribution is known, which requires knowledge of the noise parameters. Or the posterior distribution is used to marginalize over the noise parameters while evaluating the optimal statistic (Vigeland et al., 2018; Sardesai et al., 2023; Vallisneri et al., 2023), which in turn requires the same uninformative priors. Fortunately, the MVUE for an isotropic stochastic background of GWs that does not use pulsar autocorrelations is still unbiased by construction, though no longer minimum variance if suboptimal assumptions are used.
@@
In this letter, I propose a shift in perspective, advocating for the adoption of hierarchical Bayesian modeling (HBM) in PTA analysis. This approach has multiple benefits over currently used priors, chief among them being that the GW parameters will exhibit less posterior bias.
@@
In Section 2 I present a toy model that demonstrates the posterior bias of currently used uninformative priors. In Section 3 I discuss the problems of uninformative priors theoretically, and introduce Hierarchical Bayesian Modeling that should replace them. In Section 4 I give some more realistic examples, after which I briefly discuss current attempts in the literature to address posterior bias in Section 5 I. The implications of this letter on the PTA community are discussed in Section 6, after which I end with concluding remarks and an outlook in Section 7.
@@ 2.1 Toy problem
where $n_{ai}$ is the $i$-th noise process realization at detector $a$, where we use the convention that $i$ labels observations $i\in[1,N_{o}]$ and $a$ labels detectors $a\in[1,N_{p}]$. The standard deviation of the IID data in detector $a$ is given by $\sigma_{a}$. I denote a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$ as $\mathcal{N}(\mu,\sigma^{2})$.
The signal process $X_{s}$ is similar to $X_{n}$ in that it is a random process that, when considering only a single detector, produces IID data according to $\mathcal{N}(0,h^{2})$. However, when considering an array of detectors, the realizations of $X_{s}$ are correlated. The correlations that PTA projects focus on the most are so-called Hellings & Downs correlations, which represent the average correlations induced by an isotropically unpolarized ensemble of gravitational-wave sources across the sky (Hellings & Downs, 1983). These correlations $\Gamma(\zeta{ab})$ only depend on the angular separation $\zeta_{ab}$ of detector $a$ and detector $b$:
@@
The only free model parameters in the toy model are $\sigma_{a}$ and $h$. The experimental setup is defined by the number of detectors $N_{p}$ and the number of observations per detector $N_{o}$. Given the numerical values of those parameters, it is possible to generate mock realizations of $d$. In a realization of our mock universe, the values of $\sigma_{a}$ follow the distribution that represents the accuracy of a detector in our ensemble of detectors. For PTAs, this distribution represents the variation in pulsar spin stability in our Galactic population of observable millisecond pulsars. I consider four scenarios, listed in table 1, where the distribution of detector noise comes from specific distributions. These signal is of the same amplitude in expectation for all scenarios, and the expected noise amplitudes of Table 1 are tweaked so that the signal is detectable. In what follows, these different distributions will serve as examples to test the priors on the detector noise.
@@ 3 Hierarchical modeling
Pulsar Timing Array projects use data from an ensemble of pulsars, where signal and noise processes of TOAs are modeled meticulously with great precision. These pulsars come from the collection of observable Galactic MSPs, which means that the noise parameters of the MSPs come from some underlying distribution. By analyzing the data of the ensemble of pulsars jointly, it becomes possible to model this underlying distribution of model parameters. Current analyses by PTAs purposely refrain from modeling this distribution, causing a mismatch between the generative model for the data and the model used for the posterior distribution. The results in Section 2.2 show that this causes problems with inference on parameters that are covariant with the ensemble of noise parameters as a whole (i.e. the GW signal $h$). This covariance makes spin noise of particular interest. Even though for instance dispersion measure variations and pulse profile variations are also dominant sources of noise for some pulsars (e.g. Agazie et al., 2023b), it is the ensemble of spin noise parameters that is most covariant with a potential GW signal, and it is therefore the mismatch of the prior and the generative model of the spin noise that lies at the heart of the posterior biases that have been found in the literature.
In this section I propose Hierarchical Bayesian Modeling (HBM) as a solution to the situation PTAs are in, where apparent posterior bias results from using uninformative priors on the noise parameters. In some other areas of science HBMs have been adopted widely, and there is a large body of literature to draw from (for a first example in GW science, look up Adams et al., 2012). For a more general and in-depth discussion regarding multilevel and hierarchical Bayesian models, I refer to Section 3 of Loredo & Hendry (2019), who do an excellent job introducing all the subtleties of HBMs in astrophysics.
@@ 3.2 The Log-Normal HyperPrior
Note that I am free to choose the priors on the hyperparameters $\mu_{\sigma}$ and $\sigma_{\sigma}$. These particular choices are convenient, practical, and sufficient for the present discussion. Note that $\log_{10}\sigma_{\sigma}$ is the standard deviation of the prior distribution on the ensemble of noise parameters, not the prior on the individual noise parameters, denoted by $\log_{10}\sigma_{a}$. Together with Equation (1)–(4), Equation (3.2) gives a full specification of the model. Analyzing the same dataset as above with this model, I obtain the results in Figure 1 labeled as “Hierarchical”. The resulting posteriors with the Hierarchical prior are universally more consistent with injections.
Aside from checking for posterior bias by eye in Figure 1, I make use of two methods to select the best model for my datasets. The first one is Bayesian model selection. Bayes Factors can be used to select between sets of assumptions and models about the data, which means that different prior assumptions can be compared. The Bayes factors of HBM vs Log-Uniform for the “Normal”, “Box”, and “Identical” scenarios from Figure 1 are respectively $\log\mathcal{B}=33.6$, $\log\mathcal{B}=33.5$, and $\log\mathcal{B}=19.6$.
The second method to pick the best model for the datasets in this section is based on the fact that this is a toy model with mock data, so I can generate many realizations of data for which the injected parameters are known. A $p$-$p$ plot can then be used to check for consistency, which is done in Section 3.2.2.
@@ 3.2.2 Comparisons using p-p plots
where $Z=P(d\mid H)$ is the FML/evidence, $P(\theta\mid d,H)$ is the posterior distribution under model hypothesis $H$, $P(d\mid\theta,H)$ is the likelihood, and $P(\theta\mid H)$ is the prior distribution for the model parameter $\theta$. I only focus on a single model parameter here, so these distributions are one-dimensional. The cumulative posterior distribution function
@@
I now take a moment to reflect on this important result. Using this statistical test, it is possible to check whether the posterior distribution is consistent with injected values over a collection of realizations of the data generation process. But this is only true if the prior distribution used in Bayes Theorem is the same as the one of the data generation process. If a different prior is used, the posterior is not going to be statistically consistent with the injections taken over many realizations of data. If one only carries out the experiment once, on one realization of data, the posterior visually looks consistent with the injections for any reasonable prior. But, if one analyzes many realizations, they start to be sensitive to the fact that the prior used for inference should match the distribution from which the injection parameters $\theta_{\rm{inj}}$ are drawn. It is interesting to ponder the similarity of these consistency checks to what is happening in Pulsar Timing Arrays. Because an ensemble of pulsars is analyzed simultaneously, the experiment becomes sensitive to the underlying distribution that the noise parameters of pulsars were drawn from: the collection of observable Galactic MSPs.
Visually, $\xi_{r}$ and $p_{r}$ are often plotted against one another in a so called $p$-$p$ plot. These types of plots offer more information than just the value of a test. See for example Figure 13 of Lamb et al. (2023) for a detailed visual explanation of such plots. In this letter, I plot $\xi_{r}-p_{r}$ vs $p_{r}$. The mean and variance under the correct model $H$ are given by:
@@ 4 Pulsar Timing examples
In a way, this letter is done. By realizing that once an ensemble of some physical quantity is simultaneously analyzed, that the underlying distribution needs to be modeled, the straightforward solution is to introduce a HBM. Of course, a suitable distribution with suitable hyper-priors on the hyperparameters should be used, such as those of Equation (3.2). Other than that, the procedure is set.
For clarity, and to demonstrate impact, I present two more realistic examples of an HBM for pulsar timing: the First IPTA Mock Data Challenge (MDC), and an analysis of simulated data that are somewhat similar to the NANOGrav 12.5yr dataset(Alam et al., 2020). With these examples it will become clear that there is a real need for HBMs in Pulsar Timing Array science. But to analyze real-world data, I first need to introduce a slightly more flexible HBM that can account for power-law signals such as those frequently modeled in PTA projects.
@@ 4.1 Power-law PSD signals
where $a,b\in\mathbb{R}^{+}$ are hyperparameters that set the shape of the distribution. However, the values of $\log_{10}A$ and $\gamma$ parameters are probably significantly correlated under the distribution of MSPs in our collection (Lam et al., 2016), which means it is more realistic to use a multivariate distribution. Instead, I therefore propose a two-step process. First, $\gamma$ is transformed $\gamma^{\prime}=\gamma^{\prime}(\gamma)$ using an interval transform, which maps the interval $[0,7]\rightarrow\langle-\infty,\infty\rangle$. Then a multivariate Gaussian prior can be placed on $(\log_{10}A,\gamma^{\prime})$.
@@
I now define Multivariate Gaussian prior $\mathcal{N}(\mu_{h},\mathbf{\Sigma_{h}})$ ($h$ for ’hyper’) as a function of $5$ hyperparameters: two mean parameters $\mu_{h}=(\mu_{\log_{10}A},\mu_{\gamma^{\prime}})$, and three variance parameters $L_{A},L_{\gamma},L_{A,\gamma}$. The two mean parameters $\mu_{h}$ are the mean of the Multivariate Gaussian. The three $L$ parameters parameterize the non-zero elements of the Cholesky decomposition of the covariance matrix $\mathbf{\Sigma}_{h}$:
@@
Where $\mathbf{\Sigma}$ is the covariance matrix of the multivariate Gaussian. I place the following priors on these hyperparameters:
@@
Using these coordinates as the sampling parameters greatly increases the MCMC mixing rate, and allows traversing otherwise inaccessible regions of parameter space. As I noted before, if using transformations such as these, remember to include the Jacobian. The final coordinate transformation chain then becomes:

### v2

@@ Abstract
Pulsar timing array (PTA) projects have found evidence of a stochastic background of gravitational waves (GWB) using data from an ensemble of pulsars. In the literature, minimal assumptions are made about the signal and noise processes that affect data from these pulsars, such as pulsar spin noise. These assumptions are encoded as uninformative priors in Bayesian searches, though frequentist approaches make similar assumptions. Uninformative priors are not suitable for (noise) properties of pulsars in an ensemble, and they bias estimates of model parameters such as gravitational-wave signal parameters. Both frequentist and Bayesian searches are affected. In this article, more appropriate priors are proposed in the language of hierarchical Bayesian modeling, where the properties of the ensemble of pulsars are jointly described with the properties of the individual components of the ensemble. Results by PTA projects should be reevaluated using hierarchical models.
@@ 1 Introduction
Astronomy is on the cusp of a transformative era with the advent of gravitational-wave (GW) astronomy. The LIGO–Virgo collaboration has heralded this new age with observations in the higher-frequency bands (LIGO Scientific Collaboration and Virgo Collaboration et al., 2016), and now, pulsar timing arrays (PTAs) stand poised to complement these findings in the nanohertz regime with published evidence for a stochastic background of gravitational waves (GWB) (Agazie et al., 2023a; Reardon et al., 2023; Antoniadis et al., 2023; Xu et al., 2023). This advancement holds the promise of reshaping our understanding of observational cosmology, galaxy formation, and physics of the early Universe.
@@
PTAs represent a unique approach to GW detection. By precisely monitoring the arrival times of pulses from an array of millisecond pulsars (MSPs, Backer et al., 1982), these projects seek to detect the subtle perturbations caused by passing GWs (Sazhin, 1978; Detweiler, 1979; Romani, 1989; Foster & Backer, 1990). The stability of MSP emissions is key to this method; however, this stability is not without some level of irregularity. The rotational instabilities of the MSPs, often referred to as spin noise, presents a notable challenge for the PTA projects in their efforts to detect GWs (Cordes & Shannon, 2010; Lentati et al., 2016).
@@
The complexity of analyzing an ensemble of pulsars simultaneously necessitates careful statistical modeling (Taylor, 2021), particularly in the treatment of spin noise. The PTA community has approached this with Bayesian inference, where separate spin-noise components with uninformative amplitude parameters for each MSP are added to the model. While this practice seems intuitive, it has sometimes led to posterior bias in GW analyses, where we use the term “posterior bias” to describe the apparent shift of the posterior distribution with respect to the true model parameter values. Posterior bias can be a consequence of model misspecification. Multiple papers have pointed out that posterior bias occurs in contemporary analyses (Hazboun et al., 2020; Zic et al., 2022; Johnson et al., 2022), but no exact cause or satisfactory solution was given. These warning signals have been mostly ignored by subsequent searches for GWs. Those same priors are still used today, such as in recent high-profile publications (Agazie et al., 2023a; Reardon et al., 2023; Antoniadis et al., 2023; Xu et al., 2023).
@@
Even though currently used uninformative priors have been confirmed to work with injection recoveries in the literature (e.g. Lentati et al., 2013; van Haasteren et al., 2009; Ellis et al., 2013; Johnson et al., 2023, and most other methods papers on GWB detection), no rigorous statistical assessment of (posterior) bias has been possible. What has been done are studies that focus on the validity of the posterior distribution and its implementation, which can be done with $p$-$p$ plots of ensembles of mock data (e.g. Ellis et al., 2013). However, for those $p$-$p$ plots to be consistent, one needs to use the same priors for data generation as for analysis. By doing that, one glosses over a significant source of posterior bias: the use of uninformative priors on an ensemble of observable quantities that follow a different distribution in reality than assumed in the prior (Priestley, 1965). Goncharov et al. (2022) pointed out that this type of model misspecification is the cause for posterior bias, and in a way the current paper generalizes their proposed solution.
@@
Frequentist estimators are similarly affected. A minimum variance unbiased estimator (MVUE; Anholm et al., 2009; Chamberlin et al., 2015; Allen & Romano, 2023) needs to assume the data distribution is known, which requires knowledge of the noise parameters. Or the posterior distribution is used to marginalize over the noise parameters while evaluating the optimal statistic (Vigeland et al., 2018; Sardesai et al., 2023; Vallisneri et al., 2023), which in turn requires the same uninformative priors. Fortunately, the MVUE for an isotropic stochastic background of GWs that does not use pulsar autocorrelations is still unbiased by construction, though no longer minimum variance if suboptimal assumptions are used.
@@
In this letter, we propose a shift in perspective, advocating for the adoption of hierarchical Bayesian modeling (HBM) in PTA analysis. This approach has multiple benefits over currently used priors, chief among them being that the GW parameters will exhibit less posterior bias.
@@
In Section 2 we present a toy model that demonstrates the posterior bias of currently used uninformative priors. In Section 3 we discuss the problems of uninformative priors theoretically, and introduce Hierarchical Bayesian Modeling that should replace them. In Section 4 we give some more realistic examples, after which we briefly discuss current attempts in the literature to address posterior bias in Section 5 I. The implications of this letter on the PTA community are discussed in Section 6, after which we end with concluding remarks and an outlook in Section 7.
@@ 2.1 Toy problem
where $n_{ai}$ is the $i$th noise process realization at detector $a$, where we use the convention that $i$ labels observations $i\in[1,N_{o}]$ and $a$ labels detectors $a\in[1,N_{p}]$. The standard deviation of the IID data in detector $a$ is given by $\sigma_{a}$. we denote a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$ as $\mathcal{N}(\mu,\sigma^{2})$.
The signal process $X_{s}$ is similar to $X_{n}$ in that it is a random process that, when considering only a single detector, produces IID data according to $\mathcal{N}(0,h^{2})$. However, when considering an array of detectors, the realizations of $X_{s}$ are correlated. The correlations that PTA projects focus on the most are so-called Hellings & Downs correlations, which represent the average correlations induced by an isotropically unpolarized ensemble of GW sources across the sky (Hellings & Downs, 1983). These correlations $\Gamma(\zeta_{ab})$ only depend on the angular separation $\zeta_{ab}$ of detector $a$ and detector $b$:
@@
The only free model parameters in the toy model are $\sigma_{a}$ and $h$. The experimental setup is defined by the number of detectors $N_{p}$ and the number of observations per detector $N_{o}$. Given the numerical values of those parameters, it is possible to generate mock realizations of $d$. In a realization of our mock universe, the values of $\sigma_{a}$ follow the distribution that represents the accuracy of a detector in our ensemble of detectors. For PTAs, this distribution represents the variation in pulsar spin stability in our Galactic population of observable millisecond pulsars. We consider four scenarios, listed in table 1, where the distribution of detector noise comes from specific distributions. These signal is of the same amplitude in expectation for all scenarios, and the expected noise amplitudes of Table 1 are tweaked so that the signal is detectable. In what follows, these different distributions will serve as examples to test the priors on the detector noise.
@@ 3 Hierarchical modeling
PTA projects use data from an ensemble of pulsars, where signal and noise processes of TOAs are modeled meticulously with great precision. These pulsars come from the collection of observable Galactic MSPs, which means that the noise parameters of the MSPs come from some underlying distribution. By analyzing the data of the ensemble of pulsars jointly, it becomes possible to model this underlying distribution of model parameters. Current analyses by PTAs purposely refrain from modeling this distribution, causing a mismatch between the generative model for the data and the model used for the posterior distribution. The results in Section 2.2 show that this causes problems with inference on parameters that are covariant with the ensemble of noise parameters as a whole (i.e. the GW signal $h$). This covariance makes spin noise of particular interest. Even though for instance dispersion measure variations and pulse profile variations are also dominant sources of noise for some pulsars (e.g. Agazie et al., 2023b), it is the ensemble of spin noise parameters that is most covariant with a potential GW signal, and it is therefore the mismatch of the prior and the generative model of the spin noise that lies at the heart of the posterior biases that have been found in the literature.
In this section we propose Hierarchical Bayesian Modeling (HBM) as a solution to the situation PTAs are in, where apparent posterior bias results from using uninformative priors on the noise parameters. In some other areas of science HBMs have been adopted widely, and there is a large body of literature to draw from (for a first example in GW science, look up Adams et al., 2012). For a more general and in-depth discussion regarding multilevel and hierarchical Bayesian models, we refer to Section 3 of Loredo & Hendry (2019), who do an excellent job introducing all the subtleties of HBMs in astrophysics.
@@ 3.2 The Log-Normal HyperPrior
Note that we are free to choose the priors on the hyperparameters $\mu_{\sigma}$ and $\sigma_{\sigma}$. These particular choices are convenient, practical, and sufficient for the present discussion. Note that $\log_{10}\sigma_{\sigma}$ is the standard deviation of the prior distribution on the ensemble of noise parameters, not the prior on the individual noise parameters, denoted by $\log_{10}\sigma_{a}$. Together with Equation (1)–(4), Equation (3.2) gives a full specification of the model. Analyzing the same dataset as above with this model, we obtain the results in Figure 1 labeled as “Hierarchical”. The resulting posteriors with the Hierarchical prior are universally more consistent with injections.
Aside from checking for posterior bias by eye in Figure 1, we make use of two methods to select the best model for my datasets. The first one is Bayesian model selection. Bayes Factors can be used to select between sets of assumptions and models about the data, which means that different prior assumptions can be compared. The Bayes factors of HBM vs Log-Uniform for the “Normal”, “Box”, and “Identical” scenarios from Figure 1 are respectively $\log\mathcal{B}=33.6$, $\log\mathcal{B}=33.5$, and $\log\mathcal{B}=19.6$.
The second method to pick the best model for the datasets in this section is based on the fact that this is a toy model with mock data, so we can generate many realizations of data for which the injected parameters are known. A $p$-$p$ plot can then be used to check for consistency, which is done in Section 3.2.2.
@@ 3.2.2 Comparisons using p-p plots
where $Z=P(d\mid H)$ is the FML/evidence, $P(\theta\mid d,H)$ is the posterior distribution under model hypothesis $H$, $P(d\mid\theta,H)$ is the likelihood, and $P(\theta\mid H)$ is the prior distribution for the model parameter $\theta$. We only focus on a single model parameter here, so these distributions are one-dimensional. The cumulative posterior distribution function
@@
We now take a moment to reflect on this important result. Using this statistical test, it is possible to check whether the posterior distribution is consistent with injected values over a collection of realizations of the data generation process. But this is only true if the prior distribution used in Bayes Theorem is the same as the one of the data generation process. If a different prior is used, the posterior is not going to be statistically consistent with the injections taken over many realizations of data. If one only carries out the experiment once, on one realization of data, the posterior visually looks consistent with the injections for any reasonable prior. But, if one analyzes many realizations, they start to be sensitive to the fact that the prior used for inference should match the distribution from which the injection parameters $\theta_{\rm{inj}}$ are drawn. It is interesting to ponder the similarity of these consistency checks to what is happening in PTAs. Because an ensemble of pulsars is analyzed simultaneously, the experiment becomes sensitive to the underlying distribution that the noise parameters of pulsars were drawn from: the collection of observable Galactic MSPs.
Visually, $\xi_{r}$ and $p_{r}$ are often plotted against one another in a so called $p$-$p$ plot. These types of plots offer more information than just the value of a test. See for example Figure 13 of Lamb et al. (2023) for a detailed visual explanation of such plots. In this letter, we plot $\xi_{r}-p_{r}$ vs $p_{r}$. The mean and variance under the correct model $H$ are given by:
@@ 4 Pulsar Timing examples
In a way, this article is now complete. By realizing that once an ensemble of some physical quantity is simultaneously analyzed, the underlying distribution needs to be modeled, the straightforward solution is to introduce an HBM. Of course, a suitable distribution with suitable hyper-priors on the hyperparameters should be used, such as those of Equation (3.2). Other than that, the procedure is set.
For clarity, and to demonstrate impact, we present two more realistic examples of an HBM for pulsar timing: the First IPTA Mock Data Challenge (MDC), and an analysis of simulated data that are somewhat similar to the NANOGrav 12.5yr dataset(Alam et al., 2020). With these examples it will become clear that there is a real need for HBMs in PTA science. But to analyze real-world data, we first need to introduce a slightly more flexible HBM that can account for power-law signals such as those frequently modeled in PTA projects.
@@ 4.1 Power-law PSD signals
where $a,b\in\mathbb{R}^{+}$ are hyperparameters that set the shape of the distribution. However, the values of $\log_{10}A$ and $\gamma$ parameters are probably significantly correlated under the distribution of MSPs in our collection (Lam et al., 2016), which means it is more realistic to use a multivariate distribution. Instead, we therefore propose a two-step process. First, $\gamma$ is transformed $\gamma^{\prime}=\gamma^{\prime}(\gamma)$ using an interval transform, which maps the interval $[0,7]\rightarrow\langle-\infty,\infty\rangle$. Then a multivariate Gaussian prior can be placed on $(\log_{10}A,\gamma^{\prime})$.
@@
We now define Multivariate Gaussian prior $\mathcal{N}(\mu_{h},\mathbf{\Sigma_{h}})$ ($h$ for ’hyper’) as a function of $5$ hyperparameters: two mean parameters $\mu_{h}=(\mu_{\log_{10}A},\mu_{\gamma^{\prime}})$, and three variance parameters $L_{A},L_{\gamma},L_{A,\gamma}$. The two mean parameters $\mu_{h}$ are the mean of the Multivariate Gaussian. The three $L$ parameters parameterize the non-zero elements of the Cholesky decomposition of the covariance matrix $\mathbf{\Sigma}_{h}$:
@@
Where $\mathbf{\Sigma}$ is the covariance matrix of the multivariate Gaussian. We place the following priors on these hyperparameters:
@@
Using these coordinates as the sampling parameters greatly increases the MCMC mixing rate, and allows traversing otherwise inaccessible regions of parameter space. As we noted before, if using transformations such as these, remember to include the Jacobian. The final coordinate transformation chain then becomes: